OpenStudy (anonymous):

perimeter of 16 ft and area of 12ft

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

u want to find the dimensions?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if x is length and y is width then xy=12 2(x+y)=16 x+y=8 y=8-x x(8-x)=12 8x-x^2=12 x^2-8x+12=0 (x-6)(x-2)=0 x=6 or 2 y=2 0r 6

OpenStudy (anonymous):

no not that

OpenStudy (anonymous):

y?

OpenStudy (anonymous):

its a square

OpenStudy (anonymous):

ohhhhhhhhhh bad

OpenStudy (anonymous):

then x^2=4x

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

the question is we need a flower garden with a perimeter of 16feet and a area of 12 square feet

OpenStudy (anonymous):

so u should put the whole prob :s

OpenStudy (anonymous):

that is the whole prob

OpenStudy (anonymous):

its not squre garden, square feet is the unit :)

OpenStudy (anonymous):

aww u till ddnt get ur answere?

OpenStudy (anonymous):

still